OpenStudy (anonymous):

1 MEDAL AND FAN QUESTION!!!!!!!!!!!!!!!!!!!! 1. M(4, 2) is the midpoint of RS The coordinates of S are (6, 1). What are the coordinates of R? (2, 3) (5, 1.5) (8, 0) (1.5, 3) 2.Look at the following number sequence. 9, 15, 21, 27, . . . Based on the pattern, what are the next two terms? 39, 45 162, 972 33, 972 33, 39

OpenStudy (mathstudent55):

The midpoint of a segment with enpoints \( (x_1, y_1)\) and \( (x_2, y_2) \) is \( (x_m, y_m) = ( \dfrac{x_1 + x_2}{2}, \dfrac{y_1 + y_2}{2} ) \) You have \(x_m\) and \(x_1\). Solve for \(x_2\). You have \(y_m\) and \(y_1\). Solve for \(y_2\).

OpenStudy (anonymous):

ok so whats the answer

OpenStudy (anonymous):

is it 3,1.5

OpenStudy (mathstudent55):

Solve these two equations: \(4 = \dfrac{6 + x_2}{2} \) \(2 = \dfrac{1 + y_2}{2} \)

OpenStudy (mathstudent55):

For 2., look at the difference between a number and the previous number.

OpenStudy (anonymous):

i just want an answer

OpenStudy (anonymous):

i got 3,1.5

OpenStudy (mathstudent55):

That's not how OS works. We help each other solve problems, not just giving out answers. I can tell you your answer to 1. is not correct.

OpenStudy (anonymous):

1. (2,3) 2. 33, 39

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

You're welcome.